# MoPINNEnKF: Iterative Model Inference using generic-PINN-based ensemble Kalman filter

### v1

@@ Abstract
Physics-informed neural networks (PINNs) have emerged as a powerful tool for solving forward and inverse problems involving partial differential equations (PDEs) by incorporating physical laws into the training process. However, the performance of PINNs is often hindered in real-world scenarios involving noisy observational data and missing physics, particularly in inverse problems. In this work, we propose an iterative multi-objective PINN ensemble Kalman filter (MoPINNEnKF) framework that improves the robustness and accuracy of PINNs in both forward and inverse problems by using the ensemble Kalman filter and the non-dominated sorting genetic algorithm III (NSGA-III). Specifically, NSGA-III is used as a multi-objective optimizer that can generate various ensemble members of PINNs along the optimal Pareto front, while accounting the model uncertainty in the solution space. These ensemble members are then utilized within the EnKF to assimilate noisy observational data. The EnKF’s analysis is subsequently used to refine the data loss component for retraining the PINNs, thereby iteratively updating their parameters. The iterative procedure generates improved solutions to the PDEs. The proposed method is tested on two benchmark problems: the one-dimensional viscous Burgers equation and the time-fractional mixed diffusion-wave equation (TFMDWE). The numerical results show it outperforms standard PINNs in handling noisy data and missing physics.
@@ 1 Introduction
The rapid advancement of machine learning and artificial intelligence has profoundly influenced a wide range of scientific and engineering disciplines [43, 56, 30, 8, 3, 22, 33, 48, 31, 63, 41]. Among these advancements, physics-informed neural network (PINN) [3, 8, 39] has emerged as a powerful tool for solving complex partial differential equations (PDEs) by integrating physical laws directly into the learning process. PINN leverages the expressive capabilities of neural networks to approximate solutions to PDEs, offering a mesh-free and flexible alternative to traditional numerical methods. However, the effectiveness of PINN can be significantly impaired in the presence of noisy and sparse observational data (forward problem) [3, 60, 53], or missing physics, such as unknown coefficients in PDEs (inverse problem) [66, 26, 7, 21], which are common challenges in real-world applications.
@@
With perfect data, PINN is successfully used in forward and inverse problems. To train such a PINN, one needs to minimize a multi-objective loss function that includes the PDE residual (residual loss), initial conditions, boundary conditions (boundary loss), and data discrepancies (data loss) [49, 39, 55, 38, 62, 50, 59, 64]. They are widely used in many different problems. For example, Mao et al. [39] utilized PINNs to infer density, velocity, and pressure fields for the one-dimensional Euler equations based on observed density gradient data. Similarly, Rasht et al. [50] employed PINNs for full waveform inversions in seismic imaging to determine wave speed from observational data. Despite the success of PINNs, there are still several challenges. First, PINN uses the soft constraints which tends to minimize the sum of the PDE residual, boundary, and data losses with appropriate weights, however, the imbalance among different loss functions during training period may result in unusual expensive training costs [58]. This can happen when certain terms dominate or vanish prematurely, resulting in inefficient training and therefore leading to non-optimal results. Second, PINNs are usually sensitive to available noisy data, which in fact is a very common setting in real-world applications. Indeed, noisy or imperfect data can “mislead” the training and therefore generate inaccurate results with inappropriate neural network parameters; the model errors in forward problem or only missing physics in inverse problem put the problem more challenging. Third, in inverse problems where only partial physics are known, i.e., PDEs with unknown parameters, the accuracy of traditional PINNs diminishes. This is because the PDE residual loss relies on both the neural network derivatives and the unknown parameters, which makes accurate inference challenging without balanced losses in training and access to high-quality data.
@@
This paper proposes an integrated MoPINNEnKF framework that combines the Multi-objective Physics Informed Neural Network with ensemble Kalman Filter (EnKF) [15, 16, 28] to solve PDEs in both forward and inverse settings with noisy observational data. The MoPINNEnKF uses reference-point-based non-dominated sorting approach (hereby, referred to as NSGA-III) [10, 11, 12] to solve the multi-objective loss function in the original PINN. Specifically, MoPINNEnKF employs NSGA-III to generate ensemble members of PINNs within the optimal Pareto front where these ensemble members are further used as forecast model results in EnKF, together with available observation data to iteratively refines the PINNs by updating its data loss function. The MoPINNEnKF framework integrates the advantages of two methods, evolution multi-objective optimizer, i.e., NSGA-III and ensemble Kalman filter (EnKF). The former can provide more balanced and effective training for PINNs with a multi-objective loss function while the optimal solutions from NSGA-III consist of various members on the optimal Pareto front, expressing the model uncertainty. In particular, NSGA-III treats each component of PINN loss as distinct objectives and in the training process, the non-dominated sorting and crowding distance calculation methods is employed [36]. On the other hand, the later, i.e., EnKF can be used to assimilate model and observational data to find the optimal solutions in the Bayesian sense. While the original Kalman filter only handles linear systems [28], the ensemble Kalman filter (EnKF) and its variant [1, 15, 16, 40, 4, 45] extends it to a wider range of problems by using a Monte Carlo approach. Essentially, the EnKF begins with a probability distribution (represented by ensemble members in the forecast) and a likelihood function for observed data, then applies Bayes’ theorem to update this distribution (the “analysis” or posterior) once new observations are introduced. However, neural network–based PDE solvers often struggle to generate ensemble members that capture model uncertainty (model errors). The EnKF has Therefore, this MoPINNEnKF approach iterately refines the PINNs while leveraging the strengths of NSGA-III’s multi-objective optimization capabilities and EnKF’s denoising to improve the accuracy and robustness.
@@
This method can be used in both forward and inverse problem settings for the given PDEs. It is also important to note that, in the inverse problem setting for PINNs, there are two different approaches: (I) The unknown physical parameters are treated as additional independent variables and included as inputs to the neural network, which is trained over a range of parameter values [6, 5]; (II) The physical parameters are treated as trainable variables. While they do not explicitly appear in the network architecture, they affect the training process through their contribution to the loss function via the PDE residual. In this paper, we use the second approach for the inverse problem, i.e., putting unknown physics term as the trainable variables which directly contributes to the PDE residual in PINN’s loss function. In brief summary, our primary contribution includes the following:
@@
    We proposed the novel MoPINNEnKF framework that integrates Ensemble Kalman Filter (EnKF) with the NSGA-III-optimized PINN ensemble. The EnKF utilizes solutions from the Pareto front (as predictions) and assimilates observational data to update the state variables, which in turn refines the PINN training process. Experiments on both forward and inverse PDE/FPDE problems show that the framework significantly improves prediction accuracy in the presence of model imperfections and noisy data.
@@
    We demonstrate that MoPINNEnKF significantly improves prediction accuracy over traditional PINNs in scenarios with incomplete physical knowledge and noisy observations. While conventional PINNs often fail under these conditions due to loss imbalance and sensitivity to noise, our framework leverages filtered observational data via EnKF to guide the model toward the ground truth.
@@
    MoPINNEnKF also achieves strong performance in inverse PDE and FPDE problems, effectively recovering missing physical information by combining ensemble-based data assimilation with multi-objective training.
Compared to methods that rely solely on multi-objective optimization, the MoPINNEnKF framework shows substantial improvements when only noisy data are available. The proposed framework introduces a novel iterative scheme that integrates multi-objective optimization with the Kalman filter to identify optimal solutions under noise-dominated conditions—an approach not previously explored in the PINN literature. The rest of the paper is organized as follows. In Section 2, we introduce the proposed integrated model that combines NSGA-III optimizer (Section 2.2) and ensemble Kalman filter (Section 2.3) in the PINN framework (Section 2.1). Section 3 presents numerical test results for both forward and inverse problems for the proposed framework. Specifically, we test two different problems: (1). one-dimensional viscous Burgers equation in Section 3.1, and (2). time-fractional mixed diffusion-wave equations (TFMDWEs) in Section 3.2. Finally, conclusions and future research directions are discussed in Section 4.
@@ 2 General framework
In this section, we introduce the physics-informed neural network (PINN), the Kalman filter, and the Non-Dominated Sorting Genetic Algorithm-III (NSGA-III), and then describe the framework that integrates these components as illustrated in Figure 1 and Algorithm 2. The framework employs PINNs to integrate potentially incomplete physical information through partial differential equation (PDE)-based loss functions, while the Kalman filter assimilates observational data in real time to optimally estimate state variables with PINN models. Subsequently, NSGA-III is employed to optimize the PINN ensemble by treating various loss components (e.g., PDE residual, data mismatch) as separate objectives, efficiently exploring the parameter space to identify a set of non-dominated solutions (the Pareto front). This integrated approach ensures that each component informs and improves the others, leading to a comprehensive, data-driven methodology that remains faithful to the underlying physics.
@@ 2.1 Physics Informed Neural Network (PINN)
To illustrate the framework of physics informed neural networks (PINNs), we start with the general nonlinear PDE which takes the form [38, 29, 39]:
@@
Figure 1: Schematic of the proposed workflow for MoPINNEnKF framework for forward and inverse problem (parameter estimation or missing physics). (a) Firstly, the neural network takes spatiotemporal inputs $(x,t)$ and, through automatic differentiation, enforces partial differential equation (PDE) constraints, boundary and initial conditions (ICs/BCs), as well as noisy observational data via loss functions. (b) Secondly. a multi-objective loss and multiple candidate networks are then refined using a multi-objective optimizer (NSGA-III), yielding robust solutions consistent with both data and governing physics. (c)The main flow of MoPINNEnKF that uses ensemble output from NSGA-PINN as initial observation and combine
@@
Here $\{(x_{i},t_{i})\}$ are the number of points sampled at the initial/boundary locations and in the entire domain, respectively, and $\{u^{d}_{k}\}$ is a set of $u$ values that are accessible at $\{(x^{d}_{i},t^{d}_{i})\}$; $w_{res}$, $w_{ic}$, $w_{bc}$, and $w_{\text{data}}$ are the weights used to balance the interplay among the four loss terms. These weights affects both the convergence rates and the obtained minimum for the optimization problem (3) [46, 2], therefore it is crucial to choose their values. In [36], the authors integrate the NSGA-III algorithm into the optimization framework for PINNs, and we use the same strategy in this paper.
2.2 Non-Dominated Sorting Genetic Algorithm-III (NSGA-III)
Multi-objective optimization problems frequently arise in various scientific and engineering fields, including artificial intelligence [57], data mining [24], software engineering [32], scheduling [27], bioinformatics [51], and economics [44], many of which involve high-dimensional spaces. In what follows, we will introduce the nondominated sorting genetic algorithm (NSGA-III) for multi-objective optimization problems and how it is used to reduce uncertainty in PINN frameworks.
@@ 2.2.1 Multi-objective Optimization
To solve such optimization problems, one can consider population-based evolutionary algorithms, which are usually more effective than traditional mathematical programming in finding a set of solutions that balance different objective functions. However, it is becoming increasingly clear that expecting a single population-based optimization method to achieve convergence near the Pareto-optimal front and maintain a uniform distribution across the entire front in high-dimensional problems is impractical [9, 65, 19]. One remedy to tackle such an issue is to incorporate external mechanisms to support diversity maintenance, which could also alleviate the computational burden. Instead of exhaustively exploring the entire search space for Pareto-optimal solutions, the algorithm can initiate multiple predefined, targeted searches. The nondominated sorting genetic algorithm (NSGA-III), proposed by Deb and Jain [12], is based on this principle and has been shown to be an effective multi-objective optimization method within the evolutionary optimization framework. NSGA-III modifies NSGA-II [13] by incorporating a reference-point–based mechanism for many-objective problems, emphasizing population members that are non-dominated yet lie close to a set of specified reference points. NSGA-III initializes with a random population of size $N$. In each iteration, the user generates an offspring population of size $N$ with mutation and/or crossover operators. With a fixed population size, out of this total of $2N$ individuals, NSGA-III selects $N$ for the next iteration. Because non-dominated solutions are preferred, the following ranking scheme establishes the dominance relation as the principal criterion for individual survival. Individuals that are not strictly dominated by any other in the population are assigned rank $1$. Subsequent ranks are determined recursively: each unranked individual that is strictly dominated only by those with ranks $11,\ldots,k-$ 1 is assigned rank $k$. Intuitively, the lower an individual’s rank, the more interesting it is. Denote $F_{i}$ be the set of individuals with rank $i$, and let $i^{*}$ be the smallest integer such that

### v2

# iPINNER: An Iterative Physics-Informed Neural Network with Ensemble Kalman Filter

@@ Abstract
Physics-informed neural networks (PINNs) have emerged as a powerful tool for solving forward and inverse problems involving partial differential equations (PDEs) by incorporating physical laws into the training process. However, the performance of PINNs is often hindered in real-world scenarios involving noisy observational data and missing physics, particularly in inverse problems. In this work, we propose an iterative multi-objective PINN ensemble Kalman filter (iPINNER) framework that improves the robustness and accuracy of PINNs in both forward and inverse problems by using the ensemble Kalman filter and the non-dominated sorting genetic algorithm III (NSGA-III). Specifically, NSGA-III is used as a multi-objective optimizer that can generate various ensemble members of PINNs along the optimal Pareto front, while accounting the model uncertainty in the solution space. These ensemble members are then utilized within the EnKF to assimilate noisy observational data. The EnKF’s analysis is subsequently used to refine the data loss component for retraining the PINNs, thereby iteratively updating their parameters. The iterative procedure generates improved solutions to the PDEs. The proposed method is tested on two benchmark problems: the one-dimensional viscous Burgers equation and the time-fractional mixed diffusion-wave equation (TFMDWE). The numerical results show it outperforms standard PINNs in handling noisy data and missing physics.
@@ 1 Introduction
The rapid advancement of machine learning and artificial intelligence has profoundly influenced a wide range of scientific and engineering disciplines [olier2021transformational, thiyagalingam2022scientific, karpatne2018machine, cuomo2022scientific, cai2021physics, guo2023learn, lin2025energy, qi2020using, kharazmi2021identifiability, zheng2022data, mou2023combining]. Among these advancements, physics-informed neural network (PINN) [cai2021physics, cuomo2022scientific, mao2020physics] has emerged as a powerful tool for solving complex partial differential equations (PDEs) by integrating physical laws directly into the learning process. PINN leverages the expressive capabilities of neural networks to approximate solutions to PDEs, offering a mesh-free and flexible alternative to traditional numerical methods. However, the effectiveness of PINN can be significantly impaired in the presence of noisy and sparse observational data (forward problem) [cai2021physics, yang2021b, satyadharma2024assessing], or missing physics, such as unknown coefficients in PDEs (inverse problem) [zou2024correcting, jiang2021hybrid, chung2024hybrid, guo2021construct], which are common challenges in real-world applications.
@@
With perfect data, PINN is successfully used in forward and inverse problems. To train such a PINN, one needs to minimize a multi-objective loss function that includes the PDE residual (residual loss), initial conditions, boundary conditions (boundary loss), and data discrepancies (data loss) [raissi2019physics, mao2020physics, shukla2020physics, lu2021physics, zhang2022analyses, rasht2022physics, xu2023transfer, zhou2023damageIMAC, lu2025evolutionary, zhang2025constrained]. They are widely used in many different problems. For example, Mao et al. [mao2020physics] utilized PINNs to infer density, velocity, and pressure fields for the one-dimensional Euler equations based on observed density gradient data. Similarly, Rasht et al. [rasht2022physics] employed PINNs for full waveform inversions in seismic imaging to determine wave speed from observational data. Despite the success of PINNs, there are still several challenges. First, PINN uses the soft constraints which tend to minimize the sum of the PDE residual, boundary, and data losses with appropriate weights, however, the imbalance among different loss functions during training period may result in unusually expensive training costs [wong2022learning]. This can happen when certain terms dominate or vanish prematurely, resulting in inefficient training and therefore leading to non-optimal results. Second, PINNs are usually sensitive to available noisy data, which in fact is a very common setting in real-world applications. Indeed, noisy or imperfect data can “mislead” the training and therefore generate inaccurate results with inappropriate neural network parameters; the model errors in forward problem or only missing physics in inverse problem put the problem more challenging. Third, in inverse problems where only partial physics are known, i.e., PDEs with unknown parameters, the accuracy of traditional PINNs diminishes. This is because the PDE residual loss relies on both the neural network derivatives and the unknown parameters, which makes accurate inference challenging without balanced losses in training and access to high-quality data.
@@
This paper proposes an integrated iPINNER framework that combines the Iterative Physics Informed Neural Network with Ensemble Kalman Filter (EnKF) [Evensen_1996, evensen2003ensemble, KF_original] to solve PDEs in both forward and inverse settings with noisy observational data. The iPINNER uses reference-point-based non-dominated sorting approach (hereby, referred to as NSGA-III) [deb2000fast, deb2001controlled, deb2013evolutionary] to solve the multi-objective loss function in the original PINN. Specifically, iPINNER employs NSGA-III to generate ensemble members of PINNs within the optimal Pareto front where these ensemble members are further used as forecast model results in EnKF, together with available observation data to iteratively refines the PINNs by updating its data loss function. The iPINNER framework integrates the advantages of two methods, evolution multi-objective optimizer, i.e., NSGA-III and ensemble Kalman filter (EnKF). The former can provide more balanced and effective training for PINNs with a multi-objective loss function while the optimal solutions from NSGA-III consist of various members on the optimal Pareto front, expressing the model uncertainty. In particular, NSGA-III treats each component of PINN loss as distinct objectives and in the training process, the non-dominated sorting and crowding distance calculation methods is employed [lu2023nsga]. On the other hand, the latter, i.e., EnKF can be used to assimilate model and observational data to find the optimal solutions in the Bayesian sense. While the original Kalman filter only handles linear systems [KF_original], the ensemble Kalman filter (EnKF) and its variant [Anderson2001, Evensen_1996, evensen2003ensemble, mou2023efficient, chen2020predicting, popov2021multifidelity] extends it to a wider range of problems by using a Monte Carlo approach. Essentially, the EnKF begins with a probability distribution (represented by ensemble members in the forecast) and a likelihood function for observed data, then applies Bayes’ theorem to update this distribution (the “analysis” or posterior) once new observations are introduced. However, neural network–based PDE solvers often struggle to generate ensemble members that capture model uncertainty (model errors). The EnKF has therefore, this iPINNER approach iterately refines the PINNs while leveraging the strengths of NSGA-III’s multi-objective optimization capabilities and EnKF’s denoising to improve the accuracy and robustness. The proposed iPINNER framework is general and can be reformulated with different trial spaces once the PDE problem is recast as an optimization problem. In this work, we adopt a neural representation $u(x,t;\theta)$, where $\theta$ is the neural parameters, for the two reasons: (1) it can encode physical laws (e.g., PDE constraints, energy dissipation) and (2) it provides mesh-free automatic differentiation for residual evaluation.
@@
This method can be used in both forward and inverse problem settings for the given PDEs. It is also important to note that, in the inverse problem setting for PINNs, there are two different approaches: (I) The unknown physical parameters are treated as additional independent variables and included as inputs to the neural network, which is trained over a range of parameter values [chen2023reduced, chen2022autodifferentiable]; (II) The physical parameters are treated as trainable variables. While they do not explicitly appear in the network architecture, they affect the training process through their contribution to the loss function via the PDE residual. In this paper, we use the second approach for the inverse problem, i.e., putting unknown physics term as the trainable variables which directly contributes to the PDE residual in PINN’s loss function. In brief summary, our primary contribution includes the following:
@@
    We proposed the novel iPINNER framework that integrates Ensemble Kalman Filter (EnKF) with the NSGA-III-optimized PINN ensemble. The EnKF utilizes solutions from the Pareto front (as predictions) and assimilates observational data to update the state variables, which in turn refines the PINN training process. Experiments on both forward and inverse PDE/FPDE problems show that the framework significantly improves prediction accuracy in the presence of model imperfections and noisy data.
@@
    We demonstrate that iPINNER significantly improves prediction accuracy over traditional PINNs in scenarios with incomplete physical knowledge and noisy observations. While conventional PINNs often fail under these conditions due to loss imbalance and sensitivity to noise, our framework leverages filtered observational data via EnKF to guide the model toward the ground truth.
@@
    iPINNER also achieves strong performance in inverse PDE and FPDE problems, effectively recovering missing physical information by combining ensemble-based data assimilation with multi-objective training.
Compared to methods that rely solely on multi-objective optimization, the iPINNER framework shows substantial improvements when only noisy data are available. The proposed framework introduces a novel iterative scheme that integrates multi-objective optimization with the Kalman filter to identify optimal solutions under noise-dominated conditions—an approach not previously explored in the PINN literature. The rest of the paper is organized as follows. In Section 2, we introduce the proposed integrated model that combines NSGA-III optimizer (Section 2.2) and ensemble Kalman filter (Section 2.3) in the PINN framework (Section 2.1). Section 3 presents numerical test results for both forward and inverse problems for the proposed framework. Specifically, we test two different problems: (1). one-dimensional viscous Burgers equation in Section 3.1, and (2). time-fractional mixed diffusion-wave equations (TFMDWEs) in Section 3.2. Finally, conclusions and future research directions are discussed in Section 4.
@@ 2 General framework
In this section, we introduce the physics‐informed neural network (PINN), the Kalman filter, and the Non-Dominated Sorting Genetic Algorithm-III (NSGA-III), and then describe the framework that integrates these components as illustrated in Figure 1 and Algorithm 2. The framework employs PINNs to integrate potentially incomplete physical information through partial differential equation (PDE)‐based loss functions, while the Kalman filter assimilates observational data in real time to optimally estimate state variables with PINN models. Subsequently, NSGA-III is employed to optimize the PINN ensemble by treating various loss components (e.g., PDE residual, data mismatch) as separate objectives, efficiently exploring the parameter space to identify a set of non-dominated solutions (the Pareto front). This integrated approach ensures that each component informs and improves the others, leading to a comprehensive, data‐driven methodology that remains faithful to the underlying physics.
@@ 2.1 Physics Informed Neural Network (PINN)
To illustrate the framework of physics informed neural networks (PINNs), we start with the general nonlinear PDE which takes the form [lu2021physics, karniadakis2021physics, mao2020physics]:
@@
Figure 1: Schematic of the proposed workflow for iPINNER framework for forward and inverse problem (parameter estimation or missing physics). (a) Firstly, the neural network takes spatiotemporal inputs $(x,t)$ and, through automatic differentiation, enforces partial differential equation (PDE) constraints, boundary and initial conditions (ICs/BCs), as well as noisy observational data via loss functions. (b) Secondly. a multi‐objective loss and multiple candidate networks are then refined using a multi‐objective optimizer (NSGA-III), yielding robust solutions consistent with both data and governing physics. (c) The main flow of iPINNER that uses ensemble output from NSGA-PINN as initial observation and combine
@@
Here $\{(x_{i},t_{i})\}$ are the number of points sampled at the initial/boundary locations and in the entire domain, respectively, and $\{u^{d}_{k}\}$ is a set of $u$ values that are accessible at $\{(x^{d}_{i},t^{d}_{i})\}$; $w_{res}$, $w_{ic}$, $w_{bc}$, and $w_{\text{data}}$ are the weights used to balance the interplay among the four loss terms. These weights affects both the convergence rates and the obtained minimum for the optimization problem (3) [psaros2022meta, bai2023physics], therefore it is crucial to choose their values. In [lu2023nsga], the authors integrate the NSGA-III algorithm into the optimization framework for PINNs, and we use the same strategy in this paper.
2.2 Non‐Dominated Sorting Genetic Algorithm-III (NSGA-III)
Multi-objective optimization problems frequently arise in various scientific and engineering fields, including artificial intelligence [tian2021evolutionary], data mining [hong2021evolutionary], software engineering [li2019quality], scheduling [jozefowiez2008multi], bioinformatics [rockendorf2022design], and economics [park2024techno], many of which involve high-dimensional spaces. In what follows, we will introduce the nondominated sorting genetic algorithm (NSGA-III) for multi-objective optimization problems and how it is used to reduce uncertainty in PINN frameworks.
@@ 2.2.1 Multi-objective Optimization
To solve such optimization problems, one can consider population-based evolutionary algorithms, which are usually more effective than traditional mathematical programming in finding a set of solutions that balance different objective functions. However, it is becoming increasingly clear that expecting a single population-based optimization method to achieve convergence near the Pareto-optimal front and maintain a uniform distribution across the entire front in high-dimensional problems is impractical [deb2011multi, zitzler2000comparison, goldberg1989messy]. One remedy to tackle such an issue is to incorporate external mechanisms to support diversity maintenance, which could also alleviate the computational burden. Instead of exhaustively exploring the entire search space for Pareto-optimal solutions, the algorithm can initiate multiple predefined, targeted searches. The nondominated sorting genetic algorithm (NSGA-III), proposed by Deb and Jain [deb2013evolutionary], is based on this principle and has been shown to be an effective multi-objective optimization method within the evolutionary optimization framework. NSGA-III modifies NSGA-II [deb2002fast] by incorporating a reference-point–based mechanism for many-objective problems, emphasizing population members that are non-dominated yet lie close to a set of specified reference points. NSGA-III initializes with a random population of size $N$. In each iteration, the user generates an offspring population of size $N$ with mutation and/or crossover operators. With a fixed population size, out of this total of $2N$ individuals, NSGA-III selects $N$ for the next iteration. Because non-dominated solutions are preferred, the following ranking scheme establishes the dominance relation as the principal criterion for individual survival. Individuals that are not strictly dominated by any other in the population are assigned rank $1$. Subsequent ranks are determined recursively: each unranked individual that is strictly dominated only by those with ranks $1,\ldots,k-1$ is assigned rank $k$. Intuitively, the lower an individual’s rank, the more interesting it is. Denote $F_{i}$ be the set of individuals with rank $i$, and let $i^{*}$ be the smallest integer such that